The equation of a circle is 4x^2+4y^2-12x+32y=27. Find the chord of the circle that cuts the x-axis.


1
Expert's answer
2021-05-07T11:48:18-0400

"4x^2+4y^2-12x+32y=27 \\
4x^2-12x+9+4y^2+32y+64=27+9+64 \\
(2x-3)^2+(2y+8)^2=10^2 \\
O(3/2;-4) \ R=10 \\
y=0 \\ (2x-3)^2+(2*0+8)^2=10^2 \\ (2x-3)^2+64=100 \\ (2x-3)^2=36 \\ 2x-3=6 \\ 2x=9 \\x=4.5 \\ 2x-3=-6 \\ 2x=-3 \\ x=-1.5 \\ 
chord \ AB \ \frac{x-x_1}{x_2-x_1}=\frac{y-y_1}{y_2-y_1} \\
A(x_1;x_2), \ B(y_1;y_2) \\
x_1\in(-1.5;4.5) \ y_1 \in(0;1) \\
x_2 \in[-3.5;-1.5) \cup(4.5;6.5] \\
y_2\in[-9;0)"